OpenStudy (anonymous):

Which polynomial is a difference of two squares? x2 + 25 x2 − 25 x2 − 10 x2 + 10

OpenStudy (anonymous):

which one is the actual answer out of the ones i gave you????

OpenStudy (amistre64):

we are not running a free answering service ....

OpenStudy (anonymous):

you have to pay to get answers?

OpenStudy (amistre64):

the studying is free :)

OpenStudy (amistre64):

when you want to determine the difference between things you subtract and a square well .... a number of the for n^2

OpenStudy (anonymous):

so, what do you mean by that? i see ALOT of kids on here asking for the answer and they get it

OpenStudy (anonymous):

but how do you determine a difference of a perfect square, can you give me an example?

OpenStudy (amistre64):

and when i spot people breaking the rules for this site i suspend them

OpenStudy (amistre64):

perfect square are of the form: 1^2 = 1 2^2 = 4 3^2 = 9 4^2 = 16 etc ...

OpenStudy (amistre64):

so i would look for an option that involves subtraction; and perfect squares

OpenStudy (anonymous):

i dont understand the question though...Which polynomial is a difference of two squares?

OpenStudy (amistre64):

Which polynomial is a difference of two squares? Which polynomial finds the difference between two squares? Which polynomial subtracts two squares?

OpenStudy (anonymous):

so like x^2 + 10 would be a difference of perfect square cause 100 is a perfect square?

OpenStudy (amistre64):

addition is not finding the difference between 2 things

OpenStudy (anonymous):

so it would be x^2-10 :)

OpenStudy (amistre64):

close .... but 10 is not a perfect square

OpenStudy (anonymous):

o so a difference of a perfect square has to be a negative so it has to be x^2-25

OpenStudy (amistre64):

good x^2 - 5^2 is a difference between 2 squares

OpenStudy (anonymous):

thankyou! you helped alot now i can actually do my work haha

OpenStudy (amistre64):

:) youre welcome. and that is the goal of this site

OpenStudy (amistre64):

good luck

OpenStudy (anonymous):

i fanned you so i can hopefully come back to you for help :) !

OpenStudy (amistre64):

if im not too busy with my real job ... maybe

OpenStudy (anonymous):

yea thats fine! im just talking about another time if your online or something